• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 08:01:38 on June 4,2002

    Entry number 85225

    keyword=Shift Summary


    Crew on Shift: Mark Jones, Bodo Reitz, Hassan Ibrahim, Eric Voutier, ************ Jean-Sebastien Real

    0:10 Take pointing data for Q1_d50
    0:50 Start production data
    Beam is very unstable, lot of FSD trips, MCC continues looking to it and data taking
    is useless.
    1:45 MCC calls for beam tune
    2:15 Get back to production run with a stable beam at 89 uA.
    2:20 ROC14 crashes
    2:25 Back on data taking
    3:15 Things are pretty stable, MCC calls to tell that "mysteriously" the machine is
    now quite. They will try to get us back to 100 uA by steps.
    3:30 Coda crash
    3:45 Back to data taking
    4:20 Change to Q1_dfN
    7:20 No beam for 5 mn
    Beam is back and will stop at 8 for the maintenance
    Kinematics Q1_f00 will end there

    Run Summary
    ----------

    Run #..... Kin #..... Target
    1144 ....... Q1_d50 ... optics
    1145 ....... Q1_d50 ... LD2 .... Very dirty beam
    1146 ....... Q1_d50 ... LD2 .... ROC14 crash
    1147 ....... Q1_d50 ... LD2 ....
    1148 ....... Q1_d50
    1150 ....... Q1_d50 ...... last run
    1153 ....... Q1_dfN ..... cross normalization
    1155 ....... Q1_f00 ..... pointing
    1156 ....... Q1_f00 ..... production run
    1157 ....... Q1_f00 ..... production run
    1158 ....... Q1_f00 ..... production run
    1159 ....... Q1_f00 ..... production run
    1160 ....... Q1_f00 ..... production run
    1161 ....... Q1_f00 ..... production run
    1162 ....... Q1_f00 ..... production run